# HANS-Net: Hyperbolic Convolution and Adaptive Temporal Attention for Accurate and Generalizable Liver and Tumor Segmentation in CT Imaging

### v1

@@ 3.4    Adaptive Temporal Attention
To handle inter-slice dependencies across CT volumes, we introduce an Adaptive Temporal Attention (ATA) mechanism, shown in Figure 3. This ATA dynamically weighs neighboring slices based on learned temporal importance, allowing the network to enhance spatial continuity across 3D volumes while maintaining computational efficiency [15]. Equation (5) formulates the attention-enhanced output:
$\small\mathbf{F}_{\text{out}}=\mathbf{X}+\left(\mathbf{W}_{\text{temp}}\odot\text{reshape}\left(\text{softmax}\left(\frac{\mathbf{Q}\mathbf{K}^{\top}}{\sqrt{d_{k}}}\right)\cdot\mathbf{V}\right)\right)$ (5)
where $\mathbf{X}\in\mathbb{R}^{B\times C\times H\times W}$ is the input feature map, $B$ is the batch size, $C$ is the number of channels, and $H\times W$ represents the spatial resolution. The query $\mathbf{Q}$, key $\mathbf{K}$, and value $\mathbf{V}$ tensors are derived from $\mathbf{X}$ using optimized $1\times 1$ convolutions and reshaped to $\mathbb{R}^{B\times N\times d_{k}}$, where $N$ is the dot product of $H\cdot W$, and $d_{k}=C/8$ is the reduced embedding dimension. A scaled dot product attention mechanism computes spatial attention weights via $\text{softmax}(\mathbf{Q}\mathbf{K}^{\top}/\sqrt{d_{k}})$, which are then applied to $\mathbf{V}$ to generate an attention-weighted representation. This output is reshaped to the original spatial format and further modulated by a trained temporal weight vector $\mathbf{W}_{\text{temp}}\in\mathbb{R}^{B\times 1\times 1\times 1}$, derived from the temporal context using a multilayer perceptron. The result is then fused with the original input $\mathbf{X}$ through a residual connection, producing $\mathbf{F}_{\text{out}}$, which is enriched in both spatial and temporal dimensions. The output generated by this module is then fed into the Synaptic Plasticity unit.
@@ 3.5   Synaptic Plasticity Mechanism
We designed a synaptic plasticity mechanism inspired by Hebbian learning to enable biologically driven adaptive feature modulation within the framework [24]. Given an input feature tensor $\mathbf{X}$, the module first computes a channel-wise activity vector for each sample. Specifically, for each sample $b$ in batch $B$, we obtain a channel activity vector $\mathbf{a}_{b}\in\mathbb{R}^{C}$ by globally averaging the spatial responses of the feature tensor $b^{th}$. Next, for each sample, we computed a correlation matrix $\mathbf{C}_{b}\in\mathbb{R}^{C\times C}$ as the outer product of the activity vector $a_{b}$ with itself.
@@
where $\boldsymbol{\eta}$ is a trace vector; $\alpha$ is the plasticity learning rate that regulates the update process; and $\text{diag}(\mathbf{C})$ is the vector formed by the diagonal entries of $\mathbf{C}$. The updated trace vector $\boldsymbol{\eta^{\prime}}$ modulates a set of learnable synaptic scaling factors $\boldsymbol{\omega}\in\mathbb{R}^{C}$, producing an adaptive channel-wise gain. This gain adjusts the input tensor $\mathbf{X}$ element-wise. The modulated features are then linearly transformed by a learnable transformation matrix $\mathbf{T}\in\mathbb{R}^{C\times C}$ to capture cross-channel interactions before convolution. Finally, as shown in Equation (7), a convolution with kernel weights $\mathbf{W}_{\text{conv}}\in\mathbb{R}^{C_{\text{out}}\times C\times K\times K}$ refines the transformed features to produce the final output tensor $\mathbf{F}_{\text{out}}$:
@@ 3.6   Implicit Neural Representation
We developed a continuous coordinate-based neural representation module to support high-resolution segmentation predictions [14]. For each normalized spatial coordinate $\mathbf{p}_{b,n}\in\mathbb{R}^{2}$ from the $b$th image in the batch, where values lie within the range $[-1,1]^{2}$, we apply a positional encoding function $\gamma(\cdot)$ to effectively capture spatial information across multiple scales. This function encodes each coordinate dimension using sine and cosine functions with progressively increasing frequencies, as described in Equation (8):
@@ 3.7   Uncertainty Estimator
We have constructed an uncertainty-aware prediction module (see Figure 5) based on Monte Carlo dropout to estimate both predictive output and its associated confidence [25]. Given an input feature map $\mathbf{X}$, we first apply a spatial dropout with probability $p$ to simulate stochastic behavior at the test time. During training, a single forward pass is performed applying dropout followed by a lightweight prediction head $f_{\theta}(\cdot)$, consisting of two convolutional layers and a sigmoid activation, which outputs a confidence map in pixels. During inference, to quantify epistemic uncertainty, we perform stochastic forward passes $T$ by applying dropout independently at each iteration, these passes are indexed by a small variable $t\in\{1,2,\ldots,T\}$, where each $u^{(t)}$ in $f_{\theta}(\text{Dropout}(X)^{(t)})$ denotes the model prediction from the $t^{\text{th}}$ forward pass with a distinct dropout mask.
The mean prediction $\mu(\mathbf{u})$ is obtained by averaging these outputs, while the uncertainty is quantified by the variance $\sigma^{2}(\mathbf{u})$, calculated as the pixel variance across the $T$ samples. The formulations are provided in Equations (11) and (12):
@@ 4.1.1   Datasets
In our study, we used two distinct public datasets to evaluate the performance of our segmentation model. Among the datasets, the LiTS dataset has been used for training, while the 3D-IRCADb-01 data set has been used for cross-data set validation purposes. Table 1 summarizes the sample distribution.
Table 1: Distribution of samples across datasets.
Datasets # Cases Category Purposes
LiTS 130 CT Image Train, Val & Test
3D-IRCADb-01 20 CT Image Cross Dataset Val
Liver Tumor Segmentation Challenge (LiTS). The LiTS dataset [26] comprises 130 volumetric abdominal CT scans, each with a pixel resolution of $512\times 512$ and varying voxel spacings, with in-plane resolutions between 0.60 and 0.98 mm and slice thicknesses ranging from 0.45 to 5.0 mm. It includes a variety of liver-related abnormalities, such as tumors, metastases, and cysts. Tumor volumes vary from 38 mm3 to 1231 mm3, and the number of tumors per scan ranges from 0 to 12. From these scans, we extract a total of 15,654 2D tumor-containing slices from all three anatomical views (4,108 axial, 5,730 coronal, and 5,816 sagittal). These selected slices are then divided into training (10,957), validation (2,348), and test (2,349) sets, ensuring that both liver parenchyma and tumor regions are present for effective segmentation learning.
3D-IRCADb-01. The 3D-IRCADb-01 dataset [27] comprises 3D CT scans of 20 patients, featuring various cases with tumor counts ranging from 0 to 75 per scan and tumor sizes varying from 38 mm³ to 349 cm³. CT scans have slice thicknesses ranging from 0.625 mm to 2.5 mm. Despite the limited number of subjects, each scan provides rich spatial detail, high-quality expert annotations, and significant variability in tumor characteristics, making it a valuable benchmark for evaluating model generalization in cross-dataset scenarios.
@@ 4.1.2   Data Preprocessing
We preprocess the 3D medical images to standardize their format and intensity ranges. Area-based resampling is applied to CT volumes and nearest-neighbor interpolation to segmentation masks to preserve anatomical boundaries, while retaining only slices with visible tumors, including subtle lesions. Windowing is applied by clipping each voxel intensity to the range $[-200,400]$ and linearly rescaling it to $[0,1]$, enhancing contrast around soft tissues. Intensity normalization is applied after windowing by dividing the clipped value range into equal intervals across all slices. This preprocessing is consistently performed across axial, coronal, and sagittal views, with each 2D slice extracted along these axes and resized in-plane to $128\times 128$ pixels, preserving depth and ensuring uniform spatial dimensions [28].
@@ 5.1   Performance Assessment
As reported in Table 2, the HANS-Net framework achieves a Dice score of 96.67% for liver and 89.84% for tumors, demonstrating high segmentation precision across both large organ regions and smaller tumor structures. The corresponding IoU values of 93.79% (liver) and 82.39% (tumor) confirm strong spatial overlap between predictions and ground truth. Furthermore, in terms of boundary alignment, the model attains a low Average Symmetric Surface Distance (ASSD) of 0.54 mm for the liver and 0.89 mm for tumors, while the Volume Overlap Error (VOE) remains limited to 6.21% and 17.61%, respectively. These results validate that HANS-Net effectively integrates geometric and biologically inspired components to produce accurate and anatomically consistent liver tumor segmentation.
@@
Training Dynamics and Convergence. The training and validation curves for our proposed HANS-Net model demonstrate effective convergence and stable performance in liver and tumor segmentation tasks. As shown in Figure 6A, both training and validation losses decreased steadily from initial values around 0.7 to approximately 0.1, with training loss converging slightly faster than validation loss, indicating good model generalization without significant overfitting. The progression of the dice coefficient reveals distinct learning patterns between organ types. Training Dice scores for liver segmentation in Figure 6B rapidly increased from 10% to more than 90% within the first 20 epochs, demonstrating the model’s ability to quickly capture large anatomical structures.
In contrast, tumor Dice scores showed a more gradual improvement, reaching approximately 80% by epoch 80, reflecting the inherent difficulty of segmenting smaller, more heterogeneous tumor regions. Validation Dice scores in Figure 6D followed similar trends, with liver achieving over 90% and tumor reaching approximately 80%, confirming robust generalization performance. IoU metrics in Figure 6C and 6E mirrored the Dice score patterns, with liver IoU rapidly increasing to 90% during training and tumor IoU gradually improving to 75%. The faster convergence of liver metrics compared to tumor metrics highlights the model’s hierarchical learning approach, where larger anatomical structures are learned before fine-grained tumor boundaries. The ASSD values in Figure 6F decreased to minimal levels below 5mm for both organs, indicating a precise boundary delimitation. The VOE metrics in Figure 6G decreased substantially during early training, stabilizing around 0.2 for the liver and 0.3 for tumor, while the analysis in Figure 6H showed decreasing uncertainty levels for both correct and incorrect predictions, and the model became more confident in its predictions as training progressed. These comprehensive training dynamics validate that our HANS-Net framework effectively integrates hyperbolic convolutions, wavelet decomposition, and synaptic plasticity mechanisms to achieve robust liver tumor segmentation with distinct learning trajectories optimized for different anatomical structures.
@@ 5.2   Comparison With State-of-the-Art Methods
We compared the proposed HANS-Net model with several SOTA liver and tumor segmentation methods, including U-Net [29], U-Net++ [30], Attention U-Net [16], H-DenseUNet [31], and TransUNet [32]. The proposed HANS-Net framework consistently surpasses all SOTA methods in most metrics. Although TransUNet [32] achieved a marginally lower ASSD of 0.51 mm for liver, where H-DenseUNet [31] recorded a slightly lower VOE of 17.54% for tumors, HANS-Net achieved the highest Dice scores of 96.67% for liver and 89.84% for tumor, the highest IoU scores of 93.79% for liver and 82.39% for tumor, and the lowest average ASSD of 0.72 mm and VOE of 11.91%.
Table 3: Comparative performance of HANS-Net and SOTA models on the LiTS dataset for liver and tumor segmentation. Metrics include DSC (%), IoU (%), ASSD (mm), and VOE (%). The best and second-best results for each dataset are bolded and underlined, respectively.
[…]
@@
Furthermore, the mean segmentation quality between both organs is the most consistent and accurate, confirming the superiority of HANS-Net in anatomical boundary preservation and tumor detection accuracy. Table 3 presents detailed comparative performance in Dice, IoU, Average Symmetric Surface Distance (ASSD), and Volume Overlap Error (VOE), demonstrating the superiority of our HANS-Net framework in accurate, robust, and anatomically consistent liver and tumor segmentation. Figure 7 visualizes the superior segmentation capabilities of our approach compared to the SOTA methods tested.

### v2

@@ 3.4 Adaptive Temporal Attention
In volumetric data, the slice thickness and spacing can vary across scanners and acquisition protocols. In our study, all scans are first resampled following area-based resampling and nearest-neighbor interpolation during preprocessing for uniformity; and to handle inter-slice dependencies across CT volumes, we introduce an Adaptive Temporal Attention (ATA) mechanism, shown in Figure 3. This ATA dynamically weighs neighboring slices based on learned temporal importance, allowing the network to enhance spatial continuity across volumes while maintaining computational efficiency [20]. Unlike fully 3D convolutional approaches that process entire volumes simultaneously at high computational cost, our ATA mechanism operates on 2D slices while explicitly modeling inter-slice relationships through learned attention weights, thereby preserving volumetric continuity without sacrificing efficiency. The temporal attention component effectively reconstructs inter-slice dependencies by attending to contextually relevant neighboring slices during feature extraction. Equation (5) formulates the attention-enhanced output:
$\mathbf{F}_{\text{out}}=\mathbf{X}+\left(\mathbf{W}_{\text{temp}}\odot\text{reshape}\left(\text{softmax}\left(\frac{\mathbf{Q}\mathbf{K}^{\top}}{\sqrt{d_{k}}}\right)\cdot\mathbf{V}\right)\right)$ (5)
where $\mathbf{X}\in\mathbb{R}^{B\times C\times H\times W}$ is the input feature map, $B$ is the batch size, $C$ is the number of channels, and $H\times W$ represents the spatial resolution. The query $\mathbf{Q}$, key $\mathbf{K}$, and value $\mathbf{V}$ tensors are derived from $\mathbf{X}$ using optimized $1\times 1$ convolutions and reshaped to $\mathbb{R}^{B\times N\times d_{k}}$, where $N$ is the dot product of $H\cdot W$, and $d_{k}=C/8$ is the reduced embedding dimension. A scaled dot product attention mechanism computes spatial attention weights via $\text{softmax}(\mathbf{Q}\mathbf{K}^{\top}/\sqrt{d_{k}})$, which are then applied to $\mathbf{V}$ to generate an attention-weighted representation. This output is reshaped to the original spatial format and further modulated by a trained temporal weight vector $\mathbf{W}_{\text{temp}}\in\mathbb{R}^{B\times 1\times 1\times 1}$, derived from the temporal context using a multilayer perceptron. Thus, instead of assuming uniform temporal spacing, ATA dynamically adjusts the contribution of neighboring slices according to their feature correlations. This allows the attention map to dynamically adjust for non-uniform slice intervals, including irregular thickness, by modulating contextual influence according to observed spatial correlations rather than fixed slice indices. Furthermore, by extracting slices from multiple anatomical views, the ATA mechanism captures complementary spatial information from orthogonal directions, effectively approximating the comprehensive context of full 3D analysis. This multi-view approach with adaptive temporal weighting maintains robust inter-slice consistency while offering significant computational efficiency and reduced memory requirements.
The result is then fused with the original input $\mathbf{X}$ through a residual connection, producing $\mathbf{F}_{\text{out}}$, which is enriched in both spatial and temporal dimensions. The output generated by this module is then fed into the Synaptic Plasticity unit.
@@ 3.5 Synaptic Plasticity Mechanism
We designed a synaptic plasticity mechanism inspired by Hebbian learning to enable biologically driven adaptive feature modulation within the framework [32]. This mechanism is based on the fundamental principle of activity-dependent synaptic strengthening observed in neural systems, where connections between co-activated neurons are reinforced over time. While our implementation represents a computational approximation optimized for gradient-based learning rather than a direct model of neurophysiological processes, it incorporates the core concept of correlation-based weight adjustment into a differentiable framework that enhances feature adaptivity in medical image segmentation [33]. Given an input feature tensor $\mathbf{X}$, the module first computes a channel-wise activity vector for each sample. Specifically, for each sample $b$ in batch $B$, we obtain a channel activity vector $\mathbf{a}_{b}\in\mathbb{R}^{C}$ by globally averaging the spatial responses of the feature tensor $b^{th}$. Next, for each sample, we computed a correlation matrix $\mathbf{C}_{b}\in\mathbb{R}^{C\times C}$ as the outer product of the activity vector $a_{b}$ with itself.
@@
where $\boldsymbol{\eta}$ is a trace vector; $\alpha$ is the plasticity learning rate that regulates the update process; and $\text{diag}(\mathbf{C})$ is the vector formed by the diagonal entries of $\mathbf{C}$. This update mechanism mirrors the strengthening of synaptic weights in response to correlated neural activity, providing an inductive bias that encourages the network to adaptively emphasize informative feature channels during training. The updated trace vector $\boldsymbol{\eta^{\prime}}$ modulates a set of learnable synaptic scaling factors $\boldsymbol{\omega}\in\mathbb{R}^{C}$, producing an adaptive channel-wise gain. This gain adjusts the input tensor $\mathbf{X}$ element-wise. The modulated features are then linearly transformed by a learnable transformation matrix $\mathbf{T}\in\mathbb{R}^{C\times C}$ to capture cross-channel interactions before convolution. Finally, as shown in Equation (7), a convolution with kernel weights $\mathbf{W}_{\text{conv}}\in\mathbb{R}^{C_{\text{out}}\times C\times K\times K}$ refines the transformed features to produce the final output tensor $\mathbf{F}_{\text{out}}$:
@@ 3.6 Implicit Neural Representation
We developed a continuous coordinate-based neural representation module to support high-resolution segmentation predictions [19]. For each normalized spatial coordinate $\mathbf{p}_{b,n}\in\mathbb{R}^{2}$ from the $b$th image in the batch, where values lie within the range $1,1]^{2}$, we apply a positional encoding function $\gamma(\cdot)$ to effectively capture spatial information across multiple scales. This function encodes each coordinate dimension using sine and cosine functions with progressively increasing frequencies, as described in Equation (8):
@@ 3.7 Uncertainty Estimator
We have constructed an uncertainty-aware prediction module (see Figure 5) based on Monte Carlo dropout to estimate both predictive output and its associated confidence [34]. Given an input feature map $\mathbf{X}$, we first apply a spatial dropout with probability $p$ to simulate stochastic behavior at the test time. During training, a single forward pass is performed applying dropout followed by a lightweight prediction head $f_{\theta}(\cdot)$, consisting of two convolutional layers and a sigmoid activation, which outputs a confidence map in pixels. During inference, to quantify epistemic uncertainty, we perform stochastic forward passes $T$ by applying dropout independently at each iteration, these passes are indexed by a small variable $t\in\{1,2,\ldots,T\}$, where each $u^{(t)}$ in $f_{\theta}(\text{Dropout}(X)^{(t)})$ denotes the model prediction from the $t^{\text{th}}$ forward pass with a distinct dropout mask. The mean prediction $\mu(\mathbf{u})$ is obtained by averaging these outputs, while the uncertainty is quantified by the variance $\sigma^{2}(\mathbf{u})$, calculated as the pixel variance across the $T$ samples. The formulations are provided in Equations (11) and (12):
@@ 4.1.1 Datasets
In our study, we used two public datasets, including LiTS and AMOS 2022, to evaluate our model’s performance and generalization capability. It is worth noting that since the official LiTS test set (70 samples) was not publicly available, we limited our analysis to the 131 publicly released training cases to ensure full reproducibility for all researchers. We randomly split these available 131 samples into an 80:20 ratio, resulting in 105 cases for training and 26 cases for testing. To demonstrate the cross-dataset generalization and modality robustness of our model, we also evaluated our model on the AMOS 2022 dataset, which contains both CT and MRI scans with 15 annotated abdominal organs. Although the AMOS 2022 dataset has a larger number of samples, we randomly selected 70 samples (unseen by the model), containing both CT and MRI scans, for cross-dataset validation to assess the model’s robustness across different imaging modalities and anatomical variations.
Liver Tumor Segmentation Challenge (LiTS). The LiTS dataset [35] comprises 131 (publicly available) volumetric abdominal CT scans, each with a pixel resolution of $512\times 512$ and varying voxel spacings, with in-plane resolutions between 0.60 and 0.98 mm and slice thicknesses ranging from 0.45 to 5.0 mm. It includes a variety of liver-related abnormalities, such as tumors, metastases, and cysts. Tumor volumes vary from 38 mm3 to 1231 mm3, and the number of tumors per scan ranges from 0 to 12. From these scans, we extract a total of 15,654 2D tumor-containing slices from all three anatomical views (4,108 axial, 5,730 coronal, and 5,816 sagittal). These selected slices are then divided into training (10,957), validation (2,348), and test (2,349) sets, ensuring that both liver parenchyma and tumor regions are present for effective segmentation learning.
AMOS 2022. The AMOS dataset [36] is a large-scale abdominal multi-organ benchmark containing 500 CT and 100 MRI scans collected from multi-center, multi-scanner clinical environments. Each scan provides voxel-level annotations for 15 abdominal organs, including the spleen, kidneys, gallbladder, esophagus, liver, stomach, aorta, postcava, pancreas, adrenal glands, duodenum, bladder, and prostate. The dataset features multi-modality, multi-phase, and multi-disease cases, capturing a broad spectrum of imaging conditions and pathological variations.
@@ 4.1.2 Data Preprocessing
We preprocess the 3D medical images to standardize their format and intensity ranges. Area-based resampling is applied to CT volumes and nearest-neighbor interpolation to segmentation masks to preserve anatomical boundaries, while retaining only slices with visible tumors, including subtle lesions. Windowing is applied by clipping each voxel intensity to the range [200,400][-200,400] and linearly rescaling it to $[0,1]$, enhancing contrast around soft tissues. Intensity normalization is applied after windowing by dividing the clipped value range into equal intervals across all slices. This preprocessing is performed across axial, coronal, and sagittal views, with each 2D slice extracted from the original 3D volume at a resolution of $512\times 512$. Our framework processes these slices at a resized in-plane resolution of $128\times 128$ due to computational resource constraints. This resizing is applied uniformly to all slices to maintain depth and ensure uniform spatial dimensions [37].
@@ 5.1 Performance Assessment
As reported in Table 1, HANS-Net achieves a Dice score of 96.67% for liver and 89.84% for tumors, demonstrating high segmentation precision across both large organ regions and smaller tumor structures. The corresponding IoU values of 93.79% (liver) and 82.39% (tumor) confirm strong spatial overlap between predictions and ground truth. Furthermore, in terms of boundary alignment, the model attains a low Average Symmetric Surface Distance (ASSD) of 0.54 mm for the liver and 0.89 mm for tumors, while the Volume Overlap Error (VOE) remains limited to 6.21% and 17.61%, respectively. The observed performance gap between liver and tumor segmentation reflects the inherent difficulty of tumor delineation due to substantial heterogeneity in size, shape, contrast, and diffuse boundaries. The synaptic plasticity module and implicit neural representation branch specifically address these challenges by adaptively refining tumor-relevant features and modeling continuous boundaries. These results validate that HANS-Net effectively integrates geometric and biologically inspired components to produce accurate and anatomically consistent liver tumor segmentation.
@@
Training Dynamics and Convergence. The training and validation curves for our proposed HANS-Net model demonstrate effective convergence and stable performance in liver and tumor segmentation tasks. As shown in Figure 6A, both training and validation losses decreased steadily from initial values around 0.7 to approximately 0.1, with training loss converging slightly faster than validation loss, indicating good model generalization without significant overfitting. The progression of the Dice score reveals distinct learning patterns between organ types. Training Dice scores for liver segmentation in Figure 6B rapidly increased from 10% to more than 90% within the first 20 epochs, demonstrating the model’s ability to quickly capture large anatomical structures.
In contrast, tumor Dice scores showed a more gradual improvement, reaching approximately 80% by epoch 80, reflecting the inherent difficulty of segmenting smaller, more heterogeneous tumor regions. Validation Dice scores in Figure 6D followed similar trends, with liver achieving over 90% and tumor reaching approximately 80%, confirming robust generalization performance. IoU metrics in Figure 6C and 6E mirrored the Dice score patterns, with liver IoU rapidly increasing to 90% during training and tumor IoU gradually improving to 75%. The faster convergence of liver metrics compared to tumor metrics highlights the model’s hierarchical learning approach, where larger anatomical structures are learned before fine-grained tumor boundaries. The ASSD values in Figure 6F decreased to minimal levels below 5mm for both organs, indicating a precise boundary delimitation. The VOE metrics in Figure 6G decreased substantially during early training, stabilizing around 0.2 for the liver and 0.3 for the tumor, while the analysis in Figure 6H showed decreasing uncertainty levels for both correct and incorrect predictions, and the model became more confident in its predictions as training progressed. These comprehensive training dynamics validate that our HANS-Net framework effectively integrates hyperbolic convolutions, wavelet decomposition, and synaptic plasticity mechanisms to achieve robust liver tumor segmentation with distinct learning trajectories optimized for different anatomical structures.
@@ 5.2 Comparison With State-of-the-Art Methods
We compared the proposed HANS-Net model with several SOTA liver and tumor segmentation methods, including U-Net [38], U-Net++ [39], Attention U-Net [24], H-DenseUNet [40], TransUNet [41], and nnU-Net [42]. To ensure fair comparison with architectures originally designed for 3D or hybrid 2D/3D processing, we adapted them to our slice-based evaluation by processing CT slices individually through their 2D encoder-decoder pathways while retaining original components such as attention gates and dense blocks. Attention U-Net’s spatial attention was applied at each decoder level, and H-DenseUNet’s hybrid dense blocks were implemented in 2D convolutional layers with consistent hyperparameters. All baselines used the same preprocessing, multi-view slice extraction, Adam optimizer with a learning rate of 1e-4, batch size of 16, and training for 100 epochs, ensuring a consistent evaluation framework across all models. The proposed HANS-Net framework consistently surpasses all SOTA methods in most metrics. Although TransUNet [41] achieved a marginally lower ASSD of 0.51 mm for liver, where H-DenseUNet [40] recorded a slightly lower VOE of 17.54% for tumors. The nnU-Net [42], on the other hand, showed competitive liver segmentation performance and achieved a second-highest Dice score of 96.20% in the liver segmentation task [35]. Compared to the existing models, our model achieved the highest Dice scores of 96.67% for liver and 89.84% for tumor, which is +0.27%, and +2.92% Dice score improvement compared to nnU-Net and TransUNet, respectively. The model also recorded the highest IoU scores of 93.79% for liver and 82.39% for tumor, and the lowest average ASSD of 0.72 mm and VOE of 11.91%.
Table 2: Comparative performance of HANS-Net and SOTA models on the LiTS dataset for liver and tumor segmentation. Metrics include Dice (%), IoU (%), ASSD, and VOE (%). The best and second-best results for each dataset are bolded and underlined, respectively.
Model Organ Dice $\uparrow$ IoU $\uparrow$ ASSD($m$) $\downarrow$ VOE $\downarrow$
@@
Furthermore, the mean segmentation quality between both organs is the most consistent and accurate, confirming the superiority of HANS-Net in anatomical boundary preservation and tumor detection accuracy. Table 2 presents detailed comparative performance in Dice, IoU, ASSD, and VOE, demonstrating the superiority of our HANS-Net framework in accurate, robust, and anatomically consistent liver and tumor segmentation. Figure 7 visualizes the superior segmentation capabilities of our approach compared to the SOTA methods tested.